OpenStudy (anonymous):

phil wantsto buy some shoes originally priced at 34.80. there now on sale for 25% off the original price, there is a 6% sales tax on the shoes and phil has $25 about how much more money does he need to buy the shoes

OpenStudy (tasha):

one - 3 two - 3 three - 5 four - ? seven - ? twenty six - ? ? - 3 ? - 4

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

the answers chooses are 2.17, 59. 19, 33.50 and 2.67

OpenStudy (anonymous):

25% off would mean that the shoes cost 75% of the original cost. 0.75*34.80=26.10. 6% tax would make the final cost 1.06*26.10=27.67. 27.67-25= $2.67 needed to purchase shoes.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

sales tax = 34.8 (.06) \[= 2.088 \approx 2.09\] 34.80 (.25) = 8.7 34.8 - 8.7 = 26.1 + 2.09 =28.19 Phil need $3.19 more

OpenStudy (anonymous):

Is sales tax taken before or after the discount is applied? lol

OpenStudy (anonymous):

i dont know but erics right